\section*{References}
\begin{thebibliography}{9}

\bibitem {PDG}J.~Beringer \textit{et al}. (Particle Data Group),
Phys.\ Rev.\ D\textbf{86}, 010001 (2012).

\bibitem {gellmanlevy}

%\cite{Schwinger:1957em}
%\bibitem{Schwinger:1957em}
J.~S.~Schwinger,
%``A Theory Of The Fundamental Interactions,''
Annals Phys.\ \textbf{2}, 407 (1957);
%%CITATION = APNYA,2,407;%%
%\cite{GellMann:1960np}
%\bibitem{GellMann:1960np}
M.~Gell-Mann and M.~Levy,
%``The Axial Vector Current In Beta Decay,''
Nuovo Cim.\ \textbf{16}, 705 (1960);
%%CITATION = NUCIA,16,705;%%
%\cite{Weinberg:1966fm}
%\bibitem{Weinberg:1966fm}
S.~Weinberg,
%``Dynamical approach to current algebra,''
Phys.\ Rev.\ Lett.\ \textbf{18}, 188 (1967).
%%CITATION = PRLTA,18,188;%%

%\cite{Bizzarri:1970rk}


\bibitem {Bizzarri:1970}R.~Bizzarri, M.~Foster, P.~Gavillet, G.~Labrosse,
L.~Montanet, R.~Salmeron, P.~Villemoes, C.~Ghesquiere \textit{et al.},
%``Experimental results on the omega pi and pi pi systems as observed in the anti-p p annihilations at rest - anti-p p ---> omega0 pi+ pi-,''
Nucl.\ Phys.\ B \textbf{14}, 169 (1969).

\bibitem {Berkley}
%\cite{AlstonGarnjost:1971kv}
%\bibitem{AlstonGarnjost:1971kv}
M.~Alston-Garnjost, A.~Barbaro-Galtieri, S.~M.~Flatte, J.~H.~Friedman,
G.~R.~Lynch, S.~D.~Protopopescu, M.~S.~Rabin, F.~T.~Solmitz,
%``Observation Of An Anomaly In The Pi+ Pi- System At 980-mev,''
Phys.\ Lett.\ B \textbf{36}, 152 (1971);
%\cite{Flatte:1972rz}
%\bibitem{Flatte:1972rz}
S.~M.~Flatte, M.~Alston-Garnjost, A.~Barbaro-Galtieri, J.~H.~Friedman,
G.~R.~Lynch, S.~D.~Protopopescu, M.~S.~Rabin, F.~T.~Solmitz,
%``Analysis Of The Observed Anomaly In Pi Pi S Wave Scattering Near K Anti-k Threshold,''
Phys.\ Lett.\ \textbf{B38}, 232 (1972); S.~D.~Protopopescu,
M.~Alston-Garnjost, A.~Barbaro-Galtieri, S.~M.~Flatte, J.~H.~Friedman,
T.~A.~Lasinski, G.~R.~Lynch, M.~S.~Rabin \textit{et al.},
%``Pi pi Partial Wave Analysis from Reactions pi+ p ---> pi+ pi- Delta++ and pi+ p ---> K+ K- Delta++ at 7.1-GeV/c,''
Phys.\ Rev.\ D \textbf{7}, 1279 (1973).

%\cite{GarciaMartin:2011jx}
\bibitem{Kaminski} 
  R.~Garcia-Martin, R.~Kaminski, J.~R.~Pelaez and J.~Ruiz de Elvira,
  %``Precise determination of the f0(600) and f0(980) pole parameters from a dispersive data analysis,''
  Phys.\ Rev.\ Lett.\  {\bf 107}, 072001 (2011)
  [arXiv:1107.1635 [hep-ph]].
  %%CITATION = ARXIV:1107.1635;%%
  
  
\bibitem {buggf0}D.~V.~Bugg,
%``A Study in Depth of f0(1370),''
Eur.\ Phys.\ J.\ C \textbf{52}, 55 (2007) [arXiv:0706.1341 [hep-ex]].
%%CITATION = EPHJA,C52,55;%%  

\bibitem {Donald:1969}R.~A.~Donald \textit{et al.},
%``The four-prong annihilation of 1.2 gev/c anti-protons in hydrogen - (ii)
%pi-plus pi-plus pi-minus pi-minus pi-neutral final state,''
Nucl.\ Phys.\ B \textbf{11}, 551 (1969).

\bibitem {f0(1710)-1981-SLAC}C.~Edwards, R.~Partridge, C.~Peck, F.~Porter,
D.~Antreasyan, Y.~F.~Gu, W.~S.~Kollmann, M.~Richardson \textit{et al.},
%``Observation of an eta eta Resonance in J/psi Radiative Decays,''
Phys.\ Rev.\ Lett.\ \textbf{48}, 458 (1982).






  
\bibitem {f0(1710)}D.~Alde \textit{et al.} [Serpukhov-Brussels-Los
Alamos-Annecy(LAPP) Collaboration],
%``OBSERVATION OF A 1750-MeV NARROW MESON DECAYING INTO eta eta,''
Phys.\ Lett.\ B \textbf{182}, 105 (1986); D.~Alde \textit{et al.} [IFVE-Brussels-Annecy-Los Alamos Collaboration],
%``Production mechanism of the X (1740) meson,''
Phys.\ Lett.\ \textbf{B284}, 457 (1992);
T.~A.~Armstrong \textit{et al.} [E760 Collaboration],
%``Evidence for $\eta \eta$ resonances in $\bar{p}p$ annihilations at 2950 < $\sqrt{s}$ < 3620-MeV,''
Phys.\ Lett.\ B \textbf{307}, 394 (1993);
D.~V.~Bugg, I.~Scott, B.~S.~Zou,
V.~V.~Anisovich, A.~V.~Sarantsev, T.~H.~Burnett and S.~Sutlief,
%``Further Amplitude Analysis Of J / Psi $\to$ Gamma (Pi+ Pi- Pi+ Pi-),''
Phys.\ Lett.\ B \textbf{353}, 378 (1995);
%D.~Barberis \textit{et al.} [WA102 Collaboration],
%%``A Partial wave analysis of the centrally produced K+ K- and K0(S) K0(S) systems in p p interactions at 450-GeV/c and new information on the spin of the f(J)(1710),''
%Phys.\ Lett.\ B \textbf{453}, 305 (1999);
D.~Barberis \textit{et al.} [WA102 Collaboration],
%``A Coupled channel analysis of the centrally produced K+ K- and pi+ pi-
%final states in p p interactions at 450-GeV/c,''
Phys.\ Lett.\ B \textbf{462}, 462 (1999) [arXiv:hep-ex/9907055];
S.~Chekanov \textit{et al.} [ZEUS Collaboration],
%``Inclusive K0(S)K0(S) resonance production in ep collisions at HERA,''
Phys.\ Rev.\ Lett.\ \textbf{101}, 112003 (2008) [arXiv:0806.0807 [hep-ex]].

%\bibitem {Leutwyler}
%%\cite{Caprini:2005zr}
%%%\bibitem{Caprini:2005zr}
%I.~Caprini, G.~Colangelo and H.~Leutwyler,
%%``Mass and width of the lowest resonance in QCD,''
%Phys.\ Rev.\ Lett.\ \textbf{96}, 132001 (2006) [hep-ph/0512364];
%%%CITATION = HEP-PH/0512364;%%
%F.~J.~Yndurain, R.~Garcia-Martin and J.~R.~Pelaez,
%%``Experimental status of the pi pi isoscalar S wave at low energy: f(0)(600) pole and scattering length,''
%Phys.\ Rev.\ D \textbf{76}, 074034 (2007) [hep-ph/0701025];
%%%CITATION = HEP-PH/0701025;%%
%H.~Leutwyler,
%%``Model independent determination of the sigma pole,''
%AIP Conf.\ Proc.\ \textbf{1030}, 46 (2008) [arXiv:0804.3182 [hep-ph]];
%%%CITATION = APCPC,1030,46;%%
%%\bibitem {Pelaez1}
%%\cite{Yndurain:2007qm}
%%\bibitem{Yndurain:2007qm}
%R.~Kaminski, R.~Garcia-Martin, P.~Grynkiewicz and J.~R.~Pelaez,
%%``Sigma pole position and errors of a once and twice subtracted dispersive
%%analysis of pi-pi scattering data,''
%Nucl.\ Phys.\ Proc.\ Suppl.\ \textbf{186}, 318 (2009) [arXiv:0811.4510
%[hep-ph]];
%%%CITATION = NUPHZ,186,318;%%
%%\cite{GarciaMartin:2011jx}
%%\bibitem{GarciaMartin:2011jx}
%R.~Garcia-Martin, R.~Kaminski, J.~R.~Pelaez and J.~Ruiz de Elvira,
%%``Precise determination of the f0(600) and f0(980) pole parameters from a dispersive data analysis,''
%Phys.\ Rev.\ Lett.\ \textbf{107}, 072001 (2011) [arXiv:1107.1635 [hep-ph]].
%%%CITATION = ARXIV:1107.1635;%%

%\cite{Nambu:1961tp}
\bibitem{NJL} 
  Y.~Nambu and G.~Jona-Lasinio,
  %``Dynamical Model of Elementary Particles Based on an Analogy with Superconductivity. 1.,''
  Phys.\ Rev.\  {\bf 122}, 345 (1961);
  %%CITATION = PHRVA,122,345;%%
%\cite{Delbourgo:1982tv}
%\bibitem{Delbourgo:1982tv} 
  R.~Delbourgo and M.~D.~Scadron,
  %``Dynamical Symmetry Breaking And The Sigma Meson Mass In Quantum Chromodynamics,''
  Phys.\ Rev.\ Lett.\  {\bf 48}, 379 (1982);
  %%CITATION = PRLTA,48,379;%%
%\cite{Delbourgo:1993dk}
%\bibitem{Delbourgo:1993dk} 
  R.~Delbourgo and M.~D.~Scadron,
  %``Dynamical generation of the SU(2) linear sigma model,''
  Mod.\ Phys.\ Lett.\ A {\bf 10}, 251 (1995)
  [hep-ph/9910242];
  %%CITATION = HEP-PH/9910242;%%
%\cite{Delbourgo:1998kg}
%\bibitem{Delbourgo:1998kg} 
  R.~Delbourgo and M.~D.~Scadron,
  %``Dynamical generation of linear sigma model SU(3) Lagrangian and meson nonet mixing,''
  Int.\ J.\ Mod.\ Phys.\ A {\bf 13}, 657 (1998)
  [hep-ph/9807504].
  %%CITATION = HEP-PH/9807504;%%


\bibitem {Meinereferenzen}P.~Ko and S.~Rudaz,
%``Phenomenology of scalar and vector mesons in the linear sigma model,''
Phys.\ Rev.\ D \textbf{50}, 6877 (1994);
J.~R.~Pelaez,
%``Light scalars as tetraquarks or two-meson states from large N(c) and unitarized chiral perturbation theory,''
Mod.\ Phys.\ Lett.\ A \textbf{19}, 2879 (2004) [hep-ph/0411107];
D.~Parganlija, F.~Giacosa and D.~H.~Rischke,
%``How Universal is the Coupling in the Sigma Model?,''
AIP Conf.\ Proc.\ \textbf{1030}, 160 (2008) [arXiv:0804.3949 [hep-ph]];
D.~Parganlija, F.~Giacosa and D.~H.~Rischke,
%``Decay widths of resonances and pion scattering lengths in a globally
%invariant linear sigma model with vector mesons,''
PoS C\textbf{ONFINEMENT8}, 070 (2008) [arXiv:0812.2183 [hep-ph]];
D.~Parganlija, F.~Giacosa and D.~H.~Rischke,
%``Influence of Vector Mesons on the f0(600) Decay Width in a Linear Sigma
%Model with Global Chiral Invariance,''
arXiv:0911.3996 [nucl-th];
%%CITATION = IMPAE,A26,607;%%
%\cite{RuizdeElvira:2010cs}
%\bibitem{RuizdeElvira:2010cs} 
  J.~Ruiz de Elvira, J.~R.~Pelaez, M.~R.~Pennington and D.~J.~Wilson,
  %``Chiral Perturbation Theory, the ${1/N_c}$ expansion and Regge behaviour determine the structure of the lightest scalar meson,''
  Phys.\ Rev.\ D {\bf 84}, 096006 (2011)
  [arXiv:1009.6204 [hep-ph]];
  %%CITATION = ARXIV:1009.6204;%%
%\cite{Parganlija:2010ac}
%\bibitem{Parganlija:2010ac}
D.~Parganlija, F.~Giacosa, P.~Kovacs and G.~Wolf,
%``Masses of Axial-Vector Resonances in a Linear Sigma Model with Nf=3,''
AIP Conf.\ Proc.\ \textbf{1343}, 328 (2011) [arXiv:1011.6104 [hep-ph]];
%%CITATION = APCPC,1343,328;%%
%%CITATION = APPXA,3,963;%%
%\cite{Kovacs:2011eg}
%\bibitem{Kovacs:2011eg}
P.~Kovacs, G.~Wolf, F.~Giacosa and D.~Parganlija,
%``Zero temperature properties of mesons in a vector meson extended linear
%sigma model,''
arXiv:1102.4732 [hep-ph];
%%CITATION = ARXIV:1102.4732;%%
%%CITATION = APPXA,4,727;%%
%\cite{Parganlija:2011np}
%\bibitem{Parganlija:2011np}
D.~Parganlija,
%``Interpreting f0(600) and a0(980) as \bar q - q states from an Nf=3 Sigma Model with (Axial-)Vectors,''
Phys.\ Scripta T {\bf 150}, 014029 (2012)
[arXiv:1109.4331 [hep-ph]];
%%CITATION = ARXIV:1109.4331;%%
W.~Heupel, G.~Eichmann and C.~S.~Fischer,
%``Tetraquark bound states in a Bethe-Salpeter approach,''
arXiv:1206.5129 [hep-ph];
  %%CITATION = ARXIV:1208.0585;%%
%\cite{Parganlija:2012uf}
%\bibitem{Parganlija:2012uf} 
  D.~Parganlija, P.~Kovacs, G.~Wolf, F.~Giacosa and D.~H.~Rischke,
  %``Phenomenology of Axial-Vector Mesons from an Extended Linear Sigma Model,''
  arXiv:1208.2054 [hep-ph].
  %%CITATION = ARXIV:1208.2054;%%


\bibitem {Meinereferenzen1}
D.~Parganlija, F.~Giacosa, D.~H.~Rischke,
%``Vacuum Properties of Mesons in a Linear Sigma Model with Vector Mesons and Global Chiral Invariance,''
Phys.\ Rev.\ D \textbf{82}, 054024 (2010) [arXiv:1003.4934 [hep-ph]];
D.~Parganlija, F.~Giacosa and D.~H.~Rischke,
%``Structure of Scalar Mesons f(0)(600), a(0)(980), f(0)(1370) and
%a(0)(1450),''
Acta Phys.\ Polon.\ Supp.\ \textbf{3}, 963 (2010) [arXiv:1004.4817 [hep-ph]];
D.~Parganlija, F.~Giacosa, D.~H.~Rischke, P.~Kovacs and
G.~Wolf,
%``A Linear Sigma Model with Three Flavors and Vector and Axial-Vector
%Mesons,''
Int.\ J.\ Mod.\ Phys.\ A \textbf{26}, 607 (2011) [arXiv:1009.2250 [hep-ph]];
%\cite{Parganlija:2011mx}
%\bibitem{Parganlija:2011mx}
D.~Parganlija,
%``Resonances f0(1370) and f0(1710) as scalar \bar q-q states from an Nf=3
%Sigma Model with Vectors and Axial-Vectors,''
Acta Phys.\ Polon.\ Supp.\ \textbf{4}, 727 (2011) [arXiv:1105.3647 [hep-ph]].


%\cite{Parganlija:2012fy}
\bibitem{Paper3} 
  D.~Parganlija, P.~Kovacs, G.~Wolf, F.~Giacosa and D.~H.~Rischke,
  %``Meson vacuum phenomenology in a three-flavor linear sigma model with (axial-)vector mesons,''
  arXiv:1208.0585 [hep-ph].

\bibitem {f0(980)asqq}D.~Morgan,
%``Is The 0+ Nonet Respectable?,''
Phys.\ Lett.\ B \textbf{51}, 71 (1974);
%%CITATION = PHLTA,B51,71;%%
%\cite{Bramon:1980ni}
%\bibitem{Bramon:1980ni}
A.~Bramon and E.~Masso,
%``On The Quark Content Of Delta (980) And Other Scalar Mesons,''
Phys.\ Lett.\ B \textbf{93}, 65 (1980) [Erratum-ibid.\ B \textbf{107}, 455
(1981)];
%%CITATION = PHLTA,B93,65;%%
%\cite{Scadron:1982eg}
%\bibitem{Scadron:1982eg}
M.~D.~Scadron,
%``Spontaneous Breakdown And The Scalar Nonet,''
Phys.\ Rev.\ D \textbf{26}, 239 (1982);
%%CITATION = PHRVA,D26,239;%%
%\cite{vanBeveren:1986ea}
%\bibitem{vanBeveren:1986ea}
E.~van Beveren, T.~A.~Rijken, K.~Metzger, C.~Dullemond, G.~Rupp and
J.~E.~Ribeiro,
%``A low-lying scalar meson nonet in a unitarized meson model,''
Z.\ Phys.\ C \textbf{30}, 615 (1986) [arXiv:0710.4067 [hep-ph]].
%%CITATION = ZEPYA,C30,615;%%
%%\cite{Morgan:1990kw}
%%\bibitem{Morgan:1990kw}
%D.~Morgan and M.~R.~Pennington,
%%``AMPLITUDE ANALYSIS OF gamma gamma $\to$ pi pi FROM THRESHOLD TO 1.4-GeV,''
%Z.\ Phys.\ C \textbf{48}, 623 (1990);
%%%CITATION = ZEPYA,C48,623;%%
%%\cite{Morgan:1990fk}
%%\bibitem{Morgan:1990fk}
%D.~Morgan and M.~R.~Pennington,
%%``The Scalar meson enigma,''
%Nucl.\ Phys.\ Proc.\ Suppl.\ \textbf{21}, 37 (1991);
%%%CITATION = NUPHZ,21,37;%%
%%\cite{Morgan:1993td}
%%\bibitem{Morgan:1993td}
%D.~Morgan and M.~R.~Pennington,
%%``New data on the K anti-K threshold region and the nature of the f0 (S*),''
%Phys.\ Rev.\ D \textbf{48}, 1185 (1993);
%%%CITATION = PHRVA,D48,1185;%%
%%\cite{Morgan:1993rn}
%%\bibitem{Morgan:1993rn}
%D.~Morgan and M.~R.~Pennington,
%%``The Decay J / psi $\to$ phi (MM) demands the f0 (S*) be narrow,''
%Phys.\ Rev.\ D \textbf{48}, 5422 (1993);
%%%CITATION = PHRVA,D48,5422;%%
%%%CITATION = ZEPYA,C68,647;%%
%%\cite{Tornqvist:1995ay}
%%\bibitem{Tornqvist:1995ay}
%N.~A.~Tornqvist and M.~Roos,
%%``Resurrection of the Sigma Meson,''
%Phys.\ Rev.\ Lett.\ \textbf{76}, 1575 (1996) [arXiv:hep-ph/9511210];
%%%%CITATION = PRLTA,76,1575;%%
%C.~R.~Munz,
%%``Two-Photon Decays of Mesons in a Relativistic Quark Model,''
%Nucl.\ Phys.\ A \textbf{609}, 364 (1996) [arXiv:hep-ph/9601206].
%%\cite{Elias:1998bq}
%%\bibitem{Elias:1998bq}
%V.~Elias, A.~H.~Fariborz, F.~Shi and T.~G.~Steele,
%%``{QCD} sum rule consistency of lowest-lying q anti-q scalar resonances,''
%Nucl.\ Phys.\ A \textbf{633}, 279 (1998) [arXiv:hep-ph/9801415];
%%%CITATION = NUPHA,A633,279;%%
%R.~Delbourgo, D.~s.~Liu and M.~D.~Scadron,
%%``s anti-s dominance of the f0(980) meson,''
%Phys.\ Lett.\ B \textbf{446}, 332 (1999) [arXiv:hep-ph/9811474];
%%\cite{Ishida:1999qk}
%%\bibitem{Ishida:1999qk}
%M.~Ishida,
%%``Possible Classification of the Chiral Scalar sigma-Nonet,''
%Prog.\ Theor.\ Phys.\ \textbf{101}, 661 (1999) [arXiv:hep-ph/9902260];
%%%CITATION = PTPKA,101,661;%%
%%\cite{Minkowski:1998mf}
%%\bibitem{Minkowski:1998mf}
%P.~Minkowski and W.~Ochs,
%%``Identification of the glueballs and the scalar meson nonet of lowest
%%mass,''
%Eur.\ Phys.\ J.\ C \textbf{9}, 283 (1999) [arXiv:hep-ph/9811518];
%%%CITATION = EPHJA,C9,283;%%
%\cite{LucioMartinez:1999ha}
%%\bibitem{LucioMartinez:1999ha}
%J.~L.~Lucio Martinez and M.~Napsuciale,
%%``a_0(980) -> gamma gamma and f_0(980) -> gamma gamma: a consistent
%%description,''
%Phys.\ Lett.\ B \textbf{454}, 365 (1999) [arXiv:hep-ph/9903234];
%%%CITATION = PHLTA,B454,365;%%
%%\cite{Anisovich:1999aw}
%%\bibitem{Anisovich:1999aw}
%A.~V.~Anisovich, V.~V.~Anisovich, D.~V.~Bugg and V.~A.~Nikonov,
%%``Partial widths a0(980) --> gamma gamma, f0(980) --> gamma gamma and  q
%%anti-q classification of the lightest scalar mesons,''
%Phys.\ Lett.\ B \textbf{456}, 80 (1999) [arXiv:hep-ph/9903396];
%%%CITATION = PHLTA,B456,80;%%
%%\cite{vanBeveren:2000ag}
%%\bibitem{vanBeveren:2000ag}
%E.~van Beveren, G.~Rupp and M.~D.~Scadron,
%%``Why the f(0)(980) is mostly ssbar,''
%Phys.\ Lett.\ B \textbf{495}, 300 (2000) [Erratum-ibid.\ B \textbf{509}, 365
%(2001)] [arXiv:hep-ph/0009265];
%%%CITATION = PHLTA,B495,300;%%
%%\cite{Deandrea:2000yc}
%%\bibitem{Deandrea:2000yc}
%A.~Deandrea, R.~Gatto, G.~Nardulli, A.~D.~Polosa and N.~A.~Tornqvist,
%%``The s-sbar and K-Kbar nature of f_0(980) in D_s decays,''
%Phys.\ Lett.\ B \textbf{502}, 79 (2001) [arXiv:hep-ph/0012120];
%%%CITATION = PHLTA,B502,79;%%
%%\cite{DeFazio:2001uc}
%%\bibitem{DeFazio:2001uc}
%F.~De Fazio and M.~R.~Pennington,
%%``Probing the structure of $f_0(980)$ through radiative $\phi$ decays,''
%Phys.\ Lett.\ B \textbf{521}, 15 (2001) [arXiv:hep-ph/0104289];
%%%CITATION = PHLTA,B521,15;%%
%%\cite{Badalian:2002rc}
%%\bibitem{Badalian:2002rc}
%A.~M.~Badalian and B.~L.~G.~Bakker,
%%``Light-meson orbital excitations in the QCD string approach,''
%Phys.\ Rev.\ D \textbf{66}, 034025 (2002) [arXiv:hep-ph/0202246];
%%%CITATION = PHRVA,D66,034025;%%
%%\cite{Scadron:2003yg}
%%\bibitem{Scadron:2003yg}
%M.~D.~Scadron, G.~Rupp, F.~Kleefeld and E.~van Beveren,
%%``Ground-State Scalar $\bar{q}q$ Nonet: SU(3) Mass Splittings and Strong,
%%Electromagnetic, and Weak Decay Rates,''
%Phys.\ Rev.\ D \textbf{69}, 014010 (2004) [Erratum-ibid.\ D \textbf{69},
%059901 (2004)] [arXiv:hep-ph/0309109];
%%%CITATION = PHRVA,D69,014010;%%
%%\cite{Koll:2000ke}
%%\bibitem{Koll:2000ke}
%M.~Koll, R.~Ricken, D.~Merten, B.~C.~Metsch and H.~R.~Petry,
%%``A relativistic quark model for mesons with an instanton induced
%%interaction,''
%Eur.\ Phys.\ J.\ A \textbf{9}, 73 (2000) [arXiv:hep-ph/0008220].
%%%CITATION = EPHJA,A9,73;%%




\bibitem {f0(980)asq2q2}
R.~L.~Jaffe,
%``Multi-Quark Hadrons. 1. The Phenomenology Of (2 Quark 2 Anti-Quark)
%Mesons,''
Phys.\ Rev.\ D \textbf{15}, 267 (1977);
%\cite{Achasov:1980tb}
%\bibitem{Achasov:1980tb}
N.~N.~Achasov, S.~A.~Devyanin and G.~N.~Shestakov,
%``Is There A 'Signature' Of The Delta (980) Meson Four Quark Nature?,''
Phys.\ Lett.\ B \textbf{96}, 168 (1980).
%%%CITATION = PHLTA,B96,168;%%
%%\cite{Black:1998wt}
%%\bibitem{Black:1998wt}
%D.~Black, A.~H.~Fariborz, F.~Sannino and J.~Schechter,
%%``Putative Light Scalar Nonet,''
%Phys.\ Rev.\ D \textbf{59}, 074026 (1999) [arXiv:hep-ph/9808415];
%%%CITATION = PHRVA,D59,074026;%%
%T.~Teshima, I.~Kitamura and
%N.~Morisita,
%%``Mixing among light scalar mesons and L = 1 q anti-q scalar mesons,''
%J.\ Phys.\ G \textbf{28}, 1391 (2002) [arXiv:hep-ph/0105107];
%%%CITATION = JPHGB,G28,1391;%%
%F.~E.~Close and N.~A.~Tornqvist,
%%``Scalar mesons above and below 1 GeV,''
%J.\ Phys.\ G \textbf{28}, R249 (2002) [arXiv:hep-ph/0204205];
%%\cite{Pelaez:2004xp}
%%\bibitem{Pelaez:2004xp}
%J.~R.~Pelaez,
%%``On the nature of light scalar mesons from their large $N_c$ behavior,''
%Phys.\ Rev.\ Lett.\ \textbf{92}, 102001 (2004) [arXiv:hep-ph/0309292];
%.~Napsuciale and S.~Rodriguez,
%%``A chiral model for bar{q}q and bar{q}bar{q}qq$ mesons,''
%Phys.\ Rev.\ D \textbf{70}, 094043 (2004) [arXiv:hep-ph/0407037];
%J.~R.~Pelaez,
%%``Light scalars as tetraquarks or two-meson states from large Nc and
%%unitarized Chiral Perturbation Theory,''
%Mod.\ Phys.\ Lett.\ A \textbf{19}, 2879 (2004) [arXiv:hep-ph/0411107];
%%%CITATION = MPLAE,A19,2879;%%
%J.~R.~Pelaez,
%%``Light scalar and vector mesons in the large N(c) limit from unitarized
%%chiral perturbation theory,''
%AIP Conf.\ Proc.\ \textbf{814}, 670 (2006) [arXiv:hep-ph/0510118];
%%\bibitem{Pelaez:2006nj}
%J.~R.~Pelaez and G.~Rios,
%%``Nature of the f_0(600) from its N_c dependence at two loops in unitarized
%%Chiral Perturbation Theory,''
%Phys.\ Rev.\ Lett.\ \textbf{97}, 242002 (2006) [arXiv:hep-ph/0610397];
%%\bibitem{Fariborz:2007ai}
%A.~H.~Fariborz, R.~Jora and J.~Schechter,
%%``Two chiral nonet model with massless quarks,''
%Phys.\ Rev.\ D \textbf{77}, 034006 (2008) [arXiv:0707.0843 [hep-ph]];
%%%CITATION = PHRVA,D77,034006;%%
%A.~H.~Fariborz, R.~Jora and J.~Schechter,
%%``Global aspects of the scalar meson puzzle,''
%Phys.\ Rev.\ D \textbf{79}, 074014 (2009) [arXiv:0902.2825 [hep-ph]];
%%%CITATION = PHRVA,D79,074014;%%
%%\cite{Fariborz:2011es}
%%\bibitem{Fariborz:2011es}
%A.~H.~Fariborz,
%%``Probing Scalar Mesons Below and Above 1 GeV,''
%arXiv:1109.2630 [hep-ph].
%%%CITATION = ARXIV:1109.2630;%%


\bibitem {f0(980)asKK}T.~Barnes,
%``Two Photon Decays Support The (K Anti-K) Molecule Picture Of The S* (975)
%And Delta (980),''
Phys.\ Lett.\ B \textbf{165}, 434 (1985);
J.~D.~Weinstein and N.~Isgur,
%``K anti-K Molecules,''
Phys.\ Rev.\ D \textbf{41}, 2236 (1990);
%%CITATION = PHRVA,D41,2236;%%
M.~P.~Locher, V.~E.~Markushin and H.~Q.~Zheng,
%``Structure of f0(980) from a Coupled Channel Analysis of S-wave pi pi
%Scattering,''
Eur.\ Phys.\ J.\ C \textbf{4}, 317 (1998) [arXiv:hep-ph/9705230].
%%CITATION = EPHJA,C4,317;%%
%V.~E.~Markushin and M.~P.~Locher,
%%``Structure of the light scalar mesons from a coupled channel analysis of
%%the S-wave pi pi - K anti-K scattering,''
%Frascati Phys.\ Ser.\ \textbf{15}, 229 (1999) [arXiv:hep-ph/9906249];
%%%CITATION = 00309,15,229;%%
%%\cite{Baru:2003qq}
%%bibitem{Baru:2003qq}
%V.~Baru, J.~Haidenbauer, C.~Hanhart, Yu.~Kalashnikova and A.~E.~Kudryavtsev,
%%``Evidence that the a0(980) and f0(980) are not elementary particles,''
%Phys.\ Lett.\ B \textbf{586}, 53 (2004) [arXiv:hep-ph/0308129];
%%%CITATION = PHLTA,B586,53;%%
%%\cite{Bugg:2006sz}
%%\bibitem{Bugg:2006sz}
%D.~V.~Bugg,
%%``Decays of sigma, kappa, a0(980) and f0(980),''
%Eur.\ Phys.\ J.\ C \textbf{47}, 57 (2006) [arXiv:hep-ph/0603089];
%%%CITATION = EPHJA,C47,57;%%
%T.~Branz, T.~Gutsche and V.~E.~Lyubovitskij,
%%``Two-photon decay of heavy hadron molecules,''
%Phys.\ Rev.\ D \textbf{82}, 054010 (2010) [arXiv:1007.4311 [hep-ph]].
%%CITATION = PHRVA,D82,054010;%%



%\cite{Jaffe:1975fd}


\bibitem {f0(980)-as-glueball}R.~L.~Jaffe, K.~Johnson,
%``Unconventional States of Confined Quarks and Gluons,''
Phys.\ Lett.\ \textbf{B60}, 201 (1976);
%\cite{Robson:1977pm}
%\bibitem{Robson:1977pm}
D.~Robson,
%``A Basic Guide For The Glueball Spotter,''
Nucl.\ Phys.\ B \textbf{130}, 328 (1977).
%%CITATION = NUPHA,B130,328;%%


%\cite{Surovtsev:2008xr}


\bibitem {f0(980)etaeta-f0(1500)glueball-f0(1370)-f0(1710)qq}Yu.~S.~Surovtsev,
P.~Bydzovsky, R.~Kaminski and M.~Nagy,
%``Light-meson spectroscopy and combined analysis of processes with
%pseudoscalar mesons,''
Phys.\ Rev.\ D \textbf{81}, 016001 (2010) [arXiv:0811.0906 [nucl-th]].
%%CITATION = PHRVA,D81,016001;%%

%\cite{Molina:2008jw}
\bibitem{Oset} 
  R.~Molina, D.~Nicmorus and E.~Oset,
  %``The rho rho interaction in the hidden gauge formalism and the f(0)(1370) and f(2)(1270) resonances,''
  Phys.\ Rev.\ D {\bf 78}, 114018 (2008)
  [arXiv:0809.2233 [hep-ph]];
  %%CITATION = ARXIV:0809.2233;%%
%\cite{Geng:2008gx}
%\bibitem{Geng:2008gx} 
  L.~S.~Geng and E.~Oset,
  %``Vector meson-vector meson interaction in a hidden gauge unitary approach,''
  Phys.\ Rev.\ D {\bf 79}, 074009 (2009)
  [arXiv:0812.1199 [hep-ph]].
  %%CITATION = ARXIV:0812.1199;%%


\bibitem {Close:1987}F.~E.~Close,
%``Gluonic Hadrons,''
Rept.\ Prog.\ Phys.\ \textbf{51}, 833 (1988);
%%CITATION = RPPHA,51,833;%%
S.~Janowski, D.~Parganlija, F.~Giacosa and D.~H.~Rischke,
%``The Glueball in a Chiral Linear Sigma Model with Vector Mesons,''
Phys.\ Rev.\ D \textbf{84}, 054007 (2011) [arXiv:1103.3238 [hep-ph]].


\bibitem {Stani}C.~Morningstar and M.~J.~Peardon,
%``Simulating the scalar glueball on the lattice,''
AIP Conf.\ Proc.\ \textbf{688}, 220 (2004) [arXiv:nucl-th/0309068];
%%CITATION = APCPC,688,220;%%
%\cite{Loan:2005ff}
%\bibitem{Loan:2005ff}
M.~Loan, X.~Q.~Luo and Z.~H.~Luo,
%``Monte Carlo study of glueball masses in the Hamiltonian limit of SU(3)
%lattice gauge theory,''
Int.\ J.\ Mod.\ Phys.\ A \textbf{21}, 2905 (2006) [arXiv:hep-lat/0503038];
%%CITATION = IMPAE,A21,2905;%%
%\cite{Gregory:2005yr}
%\bibitem{Gregory:2005yr}
%E.~B.~Gregory, A.~C.~Irving, C.~C.~McNeile, S.~Miller and Z.~Sroczynski,
%%``Scalar glueball and meson spectroscopy in unquenched lattice QCD with
%%improved staggered quarks,''
%PoS \textbf{LAT2005}, 027 (2006) [arXiv:hep-lat/0510066];
%\cite{Chen:2005mg}
%\bibitem{Chen:2005mg}
Y.~Chen \textit{et al.},
%``Glueball spectrum and matrix elements on anisotropic lattices,''
Phys.\ Rev.\ D \textbf{73}, 014516 (2006) [arXiv:hep-lat/0510074];
%%CITATION = PHRVA,D73,014516;%%
S.~Janowski, D.~Parganlija, F.~Giacosa and D.~H.~Rischke,
%``The Glueball in a Chiral Linear Sigma Model with Vector Mesons,''
Phys.\ Rev.\ D \textbf{84}, 054007 (2011) [arXiv:1103.3238 [hep-ph]].





\bibitem {f0(1710)asglueball}E.~Klempt and A.~Zaitsev,
%``Glueballs, Hybrids, Multiquarks. Experimental facts versus QCD inspired
%concepts,''
Phys.\ Rept.\ \textbf{454}, 1 (2007) [arXiv:0708.4016 [hep-ph]];
%%CITATION = PRPLC,454,1;%%
%\cite{Albaladejo:2008qa}
%\bibitem{Albaladejo:2008qa}
M.~Albaladejo and J.~A.~Oller,
%``Identification of a Scalar Glueball,''
Phys.\ Rev.\ Lett.\ \textbf{101}, 252002 (2008) [arXiv:0801.4929 [hep-ph]].
%%CITATION = PRLTA,101,252002;%%









\bibitem {Schechter}
%\bibitem{Black:1998wt}
D.~Black, A.~H.~Fariborz, F.~Sannino and J.~Schechter,
%``Putative light scalar nonet,''
Phys.\ Rev.\ D \textbf{59}, 074026 (1999) [hep-ph/9808415];
D.~Black, A.~H.~Fariborz and J.~Schechter,
%``Mechanism for a next-to-lowest lying scalar meson nonet,''
Phys.\ Rev.\ D \textbf{61}, 074001 (2000) [hep-ph/9907516];
%%CITATION = HEP-PH 0407037;%%
%%CITATION = HEP-PH/9808415;%%
%%CITATION = HEP-PH/9907516;%%
%\cite{Lenaghan:2000ey}
%\cite{Fariborz:2005gm}
%\bibitem{Fariborz:2005gm}
A.~H.~Fariborz, R.~Jora and J.~Schechter,
%``Toy model for two chiral nonets,''
Phys.\ Rev.\ D \textbf{72}, 034001 (2005) [hep-ph/0506170];
%%CITATION = HEP-PH/0506170;%%
%\cite{Fariborz:2006ff}
F.~Giacosa,
%``Mixing of scalar tetraquark and quarkonia states in a chiral approach,''
Phys.\ Rev.\ D \textbf{75}, 054007 (2007) [arXiv:hep-ph/0611388].
%\bibitem{Fariborz:2006ff}
%A.~H.~Fariborz, R.~Jora and J.~Schechter,
%%``Model for light scalars in QCD,''
%Phys.\ Rev.\ D \textbf{76}, 014011 (2007) [hep-ph/0612200];
%%%CITATION = HEP-PH/0612200;%%
%%\cite{Fariborz:2007ai}
%%\bibitem{Fariborz:2007ai}
%A.~H.~Fariborz, R.~Jora and J.~Schechter,
%%``Two chiral nonet model with massless quarks,''
%Phys.\ Rev.\ D \textbf{77}, 034006 (2008) [arXiv:0707.0843 [hep-ph]];
%%%CITATION = ARXIV:0707.0843;%%
%%\cite{Fariborz:2009cq}
%%\bibitem{Fariborz:2009cq}
%A.~H.~Fariborz, R.~Jora and J.~Schechter,
%%``Global aspects of the scalar meson puzzle,''
%Phys.\ Rev.\ D \textbf{79}, 074014 (2009) [arXiv:0902.2825 [hep-ph]];
%%%CITATION = ARXIV:0902.2825;%%
%%\cite{Fariborz:2010ez}
%%\bibitem{Fariborz:2010ez}
%A.~H.~Fariborz, R.~Jora, J.~Schechter and M.~N.~Shahid,
%%``Are three flavors special?,''
%Phys.\ Rev.\ D \textbf{83}, 034018 (2011) [arXiv:1012.4868 [hep-ph]];
%%%CITATION = ARXIV:1012.4868;%%
%%\cite{Fariborz:2011in}
%%\bibitem{Fariborz:2011in}
%A.~H.~Fariborz, R.~Jora, J.~Schechter and M.~N.~Shahid,
%%``Chiral Nonet Mixing in pi pi Scattering,''
%Phys.\ Rev.\ D \textbf{84}, 113004 (2011) [arXiv:1106.4538 [hep-ph]];
%%\cite{Mukherjee:2012xn}
%%\bibitem{Mukherjee:2012xn}
%T.~K.~Mukherjee, M.~Huang and Q.~-S.~Yan,
%%``Low-lying Scalars in an extended Linear $\sigma$ Model,''
%arXiv:1203.5717 [hep-ph].
%%%CITATION = ARXIV:1203.5717;%%



%\bibitem{BESII} 
\bibitem {BESII2004}
M.~Ablikim \textit{et al.} [BES Collaboration],
%``Resonances in J / psi ---> phi pi+ pi- and phi K+ K-,''
Phys.\ Lett.\ B \textbf{607}, 243 (2005) [arXiv:hep-ex/0411001].
%%CITATION = PHLTA,B607,243;%%


%\cite{Lutz:2009ff}
\bibitem{PANDA} 
  M.~F.~M.~Lutz {\it et al.}  [PANDA Collaboration],
  %``Physics Performance Report for PANDA: Strong Interaction Studies with Antiprotons,''
  arXiv:0903.3905 [hep-ex].
  %%CITATION = ARXIV:0903.3905;%%



\end{thebibliography}

\documentclass[a4paper]{jpconf}
\usepackage{graphicx}
\begin{document}
\title{Scalar Mesons and FAIR}

\author{Denis Parganlija}

\address{Institute for Theoretical Physics, Vienna University of Technology, Wiedner Hauptstr.\ 8-10, A-1040 Vienna, Austria}

\ead{denisp@hep.itp.tuwien.ac.at}

\begin{abstract}
I review issues related to the scalar-meson phenomenology considered of interest for the future PANDA experiments at FAIR.
\end{abstract}

\section{Introduction}

Scalar mesons are degrees of freedom of QCD (Quantum Chromodynamics) that possess quantum numbers $J^{PC} = 0^{++}$,
where $J$ is the total spin, $P$ represents parity and $C$ is the charge 
conjugation. Mesons that possess isospin $I=0$ are referred to as isoscalars. Scalars with $I=0$ represent QCD degrees of freedom with some 
peculiar features: they are defined as carrying quantum numbers
of vacuum but are 
identifiable as resonances; as mesons, they should be of $\bar q q$ (quarkonium) structure but are too many to simply 
represent $\bar q q$ objects; and they are strongly mixed with each other so that disentangling them is in the majority of 
cases a highly non-trival task.
However, they are also of extreme importance for QCD since their vacuum expectation values represent the
cause of the spontaneous breaking of the chiral symmetry in QCD, and this "Higgs mechanism" is in turn expected to generate 
Goldstone bosons of QCD: pions, kaons and others. Thus understanding isoscalar mesons
is a fundamental question of QCD, and one that has by no means been closed after decades of research and debate.
\subsection{Scalar Mesons and Experiment}

Listings of the Particle Data Group (PDG) cite the existence of 
five $IJ^{PC} = 00^{++}$ states in the energy region up to 1.8 GeV: $f_0(500)$ or $\sigma$, $f_0(980)$, $f_0(1370)$, $f_0(1500)$ and $f_0(1710)$ \cite{PDG}. Let us review 
basic experimental data regarding these resonances.
\begin{itemize}

\item Historically, debates regarding scalar mesons have been mostly focused on the $f_0(500)$ resonance (also known as 
$\sigma$). The latter
is characterised by a decay width that is of the same magnitude as the resonance mass -- according to PDG, it has a mass of (400-500) MeV and 
a decay width of (400-700) MeV. The 
existence of this state was suggested in linear sigma models long ago: the state was introduced
theoretically as the putative chiral partner of the pion \cite{gellmanlevy}. Since the pion is a 
well-established quarkonium state, then the chiral-partner state of the pion had to be of the same structure. Additionally, the pion is a
pseudo-Goldstone bosone -- its mass is the lowest in the meson spectrum. 
Therefore, naive expectations were that ({\it i}) the chiral partner of the pion would be the lightest scalar state with $I=0$, 
i.e., the $\sigma$ meson, and consequently ({\it ii}) that the $\sigma$ meson was a $\bar q q$ state. However, currently it is by no means clear that 
$\sigma$ is truly a $\bar q q$ state, implying that its interpretation as the chiral partner of the pion is in doubt.
We will come back to this issue later in this manuscript.

\item First hints about the existence of the $f_0(980)$ resonance were discovered in analyses of ${\pi\pi}$ scattering data
found to exhibit a rapid drop in the cross-section in the energy region
between $950$ MeV and $980$ MeV (i.e., close to the $KK$ threshold) in Saclay data on antiproton-proton 
\cite{Bizzarri:1970} and Berkley data on pion-proton collisions \cite{Berkley}. The most precise modern-data analyses result 
in pole mass and pole decay widths values in the close vicinity of the mentioned energy region -- see, e.g., Ref.\ \cite{Kaminski}
where the pole mass was determined as $m_{f_0(980)} = (996 \pm 7)$ MeV and the pole decay width as $\Gamma_{f_0(980)} =
50^{+20}_{-12}$ MeV.

\item The $f_0(1370)$ resonance is a broad enhancemenet in the $\pi\pi$, $4\pi$ and, to a lesser extent, $KK$ and $\eta \eta$ 
channels in the energy region of approximately 1.3 GeV. It is characterised by emergence of two peaks, respectively in the 
$2\pi$ and $4\pi$ decay channels \cite{buggf0}, the disentaglement of which requires a simultaneous analysis of $\pi\pi 
\rightarrow \pi\pi$ and $\pi \pi \rightarrow 4\pi$ scattering data. The analysis of Ref.\ \cite{buggf0} demonstrates that the 
$2\pi$ 
decay channel is dominant in the energy region up to 1.35 GeV whereas the energy region thereafter is dominated
by the opening of the $\rho \rho$ threshold leading to a strong $4\pi$ decay pattern.
Note that $f_0(1370)$ represents a single resonance despite possessing two peaks since a combined dispersive 
analysis of both $2\pi$ and $4\pi$ channels yields only one pole. 
Historically, early hints of the $f_0(1370)$ existence were observed in $\bar{p} p$ collisions at CERN in 1969 with a 
possible $\rho\rho$ enhancement claimed at $1.4$ GeV \cite{Donald:1969}. Current edition of the Particle Data Group listings
accumulates all available data on $f_{0}(1370)$ estimating an interval for, rather than stating exact values of, its mass and 
decay width: $m_{f_{0}(1370)}=(1200$-$1500)$ MeV and $\Gamma_{f_{0}(1370)}=(200$-$500)$
MeV \cite{PDG}.

\item The discovery of the $f_{0}(1500)$ resonance originated in search for the
scalar glueball state. This resonance is found mostly in pion final states
from nucleon-nucleon (or antinucleon-nucleon) and pion-nucleon scattering
processes. If such processes produce four pions, then $f_{0}(1500)$ is
reconstructed from $\rho\rho$ final states in the $2(\pi^{+}\pi^{-})$ channel
and from $\sigma\sigma$ final states in the $2(\pi^{+}\pi^{-})$ or $2(\pi
^{0}\pi^{0})$ channels. The resonance is therefore at least partly
reconstructed in channels containing a double Pomeron exchange rendering the
state a glueball candidate \cite{Close:1987}.
The PDG cites a world-average mass $m_{f_{0}(1500)}=(1505\pm6)$ MeV and decay
width $\Gamma_{f_{0}(1710)}=(109\pm7)$ MeV \cite{PDG}.

\item The $f_0(1710)$ resonance is characterised by a predominant $KK$ decay channel, marking a clear point of 
distinction between this resonance and the above mentioned ones.
The earliest evidence for the $f_{0}(1710)$ resonance was obtained from the
decay $J/\psi\rightarrow\gamma\eta\eta$ at the SLAC Crystal Ball detector from
$e^{+}e^{-}$ annihilation \cite{f0(1710)-1981-SLAC}; a
resonance with a mass of $(1640\pm50)$ MeV, a decay width of $220_{-70}^{+100}$ MeV
and the charge conjugation
$C=+1$ was found (the latter was because the resonance had been produced in a radiative $J/\psi$ decay but, interestingly, no 
final determination of the total spin and parity was possible from these first data).
Subsequent analyses, such as those in Ref.\ \cite{f0(1710)}, determined the resonance to be of $J^{P}=0^{+}$ nature.
The PDG cites a world-average mass $m_{f_{0}(1710)}=(1720\pm6)$ MeV and a decay
width $\Gamma_{f_{0}(1710)}=(135\pm8)$ MeV \cite{PDG}.
\end{itemize}

\subsection{Scalar Mesons and Theory}

Assuming a $\bar{q} q$ structure for mesons (where $q$ denotes a constituent quark) it is possible to construct two $IJ^{PC} 
= 00^{++}$ states for three quark flavours
(up $u$, down $d$ and strange $s$) once the approximate isospin symmetry in the nonstrange-quark sector is considered ($u = d$):
$
f_0^N \equiv (\bar{u}u+\bar{d}d)/\sqrt{2}$ and $f_0^S \equiv \bar{s}s$. The latter are per construction pure states. They can obviously mix by virtue of carrying the same quantum numbers; the 
mixed states are expected to be those
appearing in the physical spectrum. Clearly, two pure states will produce the same number of mixed states -- thus only two of the five previously discussed physical states
can conceivably be described in this approach. Hence the natural questions are

\begin{itemize}

\item Where in the physical spectrum are isoscalar $\bar{q} q$ states located?

\item What structure do the remaining states possess? 

\end{itemize}

Answers to these questions are complicated by the already mentioned fact that states with the same $IJ^{PC}$ quantum
numbers in general, and isoscalar mesons in particular, are expected to mix. Thus an isoscalar (or indeed any other) meson need not necessarily be of
$\bar q q$ structure -- it can also possess admixtures (or even be predominantly composed) of tetraquark ($\bar{q}\bar{q}qq$),
glueball (bound states of QCD gauge bosons -- gluons) or molecular (e.g., $\pi\pi$ or $KK$) contributions. For the cases of the above-stated isoscalar mesons we then note the following:

\begin{itemize}
\item As indicated above, the $f_0(500)$ resonance is usually considered as the natural option for the chiral partner of the pion, implying that $f_0(500)$
possesses the $\bar{q} q$ structure
as suggested, e.g., by Ref.\ \cite{NJL}. However, a $\bar
{q}q$ scalar state possesses the intrinsic angular momentum $L=1$ as well as
the relative spin of the quarks $S=1$. For this reason one could also easily
expect the scalar $\bar q q$ state to be in the region above $1$ GeV.
This would imply that ({\it i}) isoscalars above 1 GeV possess the $\bar q q$ structure and ({\it ii}) $f_0(500)$ represents a tetraquark state or a $\pi\pi$ 
bound state. For details on both statements see, e.g., Refs.\ \cite{Meinereferenzen, Meinereferenzen1, Paper3}.

\item As already discussed, the $f_0(980)$ resonance is close to the kaon-kaon threshold rendering a structure analysis for this state rather difficult. The resonance may be interpreted as a quarkonium
\cite{f0(980)asqq}, as a $\bar{q}^{2}q^{2}$ state
\cite{f0(980)asq2q2}%
, as a $KK$ bound state
\cite{f0(980)asKK}%
, as a glueball \cite{f0(980)-as-glueball} or even as an $\eta\eta$ bound
state \cite{f0(980)etaeta-f0(1500)glueball-f0(1370)-f0(1710)qq}.

\item The $f_0(1370)$ resonance appears to be a suitable candidate for a non-strange $\bar{q} q$ state \cite{Meinereferenzen1,Paper3} or a dynamically generated state
\cite{Oset}.

\item The $f_0(1500)$ resonance seems to represent a (predominant) glueball state \cite{Close:1987, Stani} although
claims have also been made that, contrarily, $f_0(1710)$ is 
of predominantly glueball nature \cite{f0(1710)asglueball}. Alternatively, $f_0(1710)$ can also be interpreted as a 
predominantly $\bar s s$ state
\cite{Paper3}.

\end{itemize}

Hence the current state of knowledge allows us to only claim with certitude that all the above mentioned states mix \cite{Schechter}; some of them even overlap due to large decay widths \cite{PDG}.
The situation is thus rather complicated -- and it is enhanced even further by the possible existence of a sixth isoscalar state, located very proximal to the set
of the already mentioned isoscalars: the $f_0(1790)$ resonance.

\section{The \boldmath $f_0(1790)$ Resonance and FAIR}

The existence of the putative new $f_0(1790)$ resonance has been claimed by the BES II Collaboration in 2004 \cite{BESII2004}. The mass and decay width of the 
resonance were determined by the Collaboration as $m_{f_{0}(1790)}=1790_{-30}^{+40}$ and $\Gamma_{f_{0}(1790)}=270_{-30}^{+60}$ MeV. The stated large decay width implies a strong overlap
of this resonance with $f_0(1710)$; however, there is a clear point of distinction between the two resonances: $f_0(1710)$ is reconstructed predominantly in the kaon
decay channels whereas $f_0(1790)$ is reconstructed predominantly in the pion decay channels.\\
There are four basic production mechanisms for $f_0(1710)$ and $f_0(1790)$ via $J/\psi$ decays \cite{BESII2004}:

\begin{itemize}

\item ({\it i}) $J/\psi\rightarrow\varphi K^+ K^-$, 

\item ({\it ii}) $J/\psi\rightarrow\varphi \pi^{+}\pi^{-}$,

\item ({\it iii}) $J/\psi\rightarrow\omega K^{+}K^{-}$, 

\item ({\it iv}) $J/\psi\rightarrow\omega \pi^{+}\pi^{-}$.

\end{itemize}

Reactions ({\it i}) and ({\it iii}) allow for reconstruction of $f_{0}(1710)$ \cite{PDG}
whereas reactions ({\it ii}) and ({\it iv}) allow for reconstruction of $f_{0}(1790)$.
Importantly, assuming $f_{0}(1710)$ and $f_{0}(1790)$ to be the same resonance leads to a contradiction: 
such a resonance would have to possess a pion-to-kaon-decay ratio of $1.82 \pm 0.33$ according to reactions
({\it i}) and ({\it ii}) and, simultaneously, the pion-to-kaon-decay ratio $< 0.11$ according to reactions
({\it iii}) and ({\it iv}) \cite{BESII2004}. There can be no one resonance with these simultaneous features -- thus
BES II data suggest $f_{0}(1710)$ and $f_{0}(1790)$ to represent two distinct resonances.\\

Let us finally discuss the relevance of the deliberations in this paper for FAIR. The proposed PANDA experiments at FAIR will study
interactions between antiprotons and fixed-target protons and nuclei in the momentum range of 1.5-15 GeV. A stated goal of the PANDA Collaboration is the
exploration of light-hadron spectroscopy, including search for glueballs and multiquark states \cite{PANDA}. In this regard I 
would like to suggest the following courses of action in the 
particular case of possible mixing scenarios in the scalar sector:

\begin{itemize}
\item Any viable search for members of a glueball spectrum must include the ground state as otherwise the spectrum would be incomplete.
The PANDA Collaboration is in a unique position to greatly advance our search for glueballs since ({\it i}) the Collaboration can build upon extensive experimental work
performed in the last decades and ({\it ii}) the technical specifications of the PANDA detector allow for a search for glueballs with unprecedented accuracy in
comparison with similar undertakings in the past.\\
However, the glueball state will inevitably overlap with other states that possess the same quantum numbers (and may be of $\bar q q$, $\bar{q}\bar{q}qq$ or molecular
structure) and the disentanglement of various signals from each of the mentioned resonances will represent a formidable task.\\

\item Current theoretical results \cite{Stani,f0(1710)asglueball} suggest the scalar glueball to be positioned at (1.5-1.7) GeV, i.e., in the vicinity of the 
established resonances $f_0(1370)$, $f_0(1500)$ and $f_0(1710)$ as well as the newly proposed $f_0(1790)$ state. Consequently, a viable search for the scalar glueball
will have to consider overlap of the glueball not only with $f_0(1370)$, $f_0(1500)$ and $f_0(1710)$ \emph{but also} with $f_0(1790)$. In other words:
a viable search for the glueball ground state at PANDA must consider the possibility of the existence of $f_0(1790)$.
If $f_0(1790)$ is not found at PANDA, then BES results may be put in doubt; but if $f_0(1790)$ is ascertained then
its signal will have to be disentangled from the glueball signal in order for the glueball search to be successful, 
making the search even more challenging and thus even more interesting.

\end{itemize}

\section{Conclusions}

Scalar mesons represent an interesting challenge for the future PANDA experiments at FAIR: the planned search for the 
glueball state in the
$IJ^{PC} = 00^{++}$ channel will have to consider mixing/overlap between non-strange $\bar q q$, strange $\bar q q$, non-strange and strange $\bar{q}\bar{q}qq$
as well as possible molecular-type states. Additionally, obtaining a clear signal for the glueball will inevitably require the confirmation (or negation)
of the existence of a putative new isoscalar state referred to as $f_0(1790)$ by the BES II Collaboration \cite{BESII2004}. If $f_0(1790)$ exists, it will most certainly overlap with the glueball. Thus
a viable search for a glueball state simultaneously implies a search for $f_0(1790)$: ignoring $f_0(1790)$ will strongly distort
any signal for the glueball. For this reason, the search for a glueball at PANDA is more than a search for merely one state -- it is a search where several very
close-by states have to be considered simultaneously.

\end{document}